# Optimal Virtual Model Control for Robotics: Design and Tuning of Passivity-Based Controllers

### v2

@@ Abstract
Index Terms:
Compliance and Impedance control, Optimization and Optimal Control, Motion Control of Manipulators, Surgical Robotics: Laparoscopy
@@ I Introduction
Passivity-based control methods are crucial for designing of stable and reliable robot controllers [1, 2, 3, 4, 5, 6, 7]. Passivity-based approaches take into account the energy of the robot-controller interconnection (energy shaping) and regulate its dissipation (damping injection) to guarantee stability. The goal is a passive controlled robot that is stable when interacting with other passive devices. The passivity theorem [8, 9] guarantees stability even if both the robot model and the environment are uncertain, as long as they are passive.
A classical example is the ubiquitous proportional derivative (PD) control plus gravity compensation, in joint or end-effector space [10, 11]. The proportional action has the mechanical interpretation of a spring. Combined with gravity compensation, it shapes the potential energy of the controlled robot. The derivative action has the mechanical interpretation of a damper, which has the effect of slowing down the robot motion towards the minimum of the shaped potential energy (stabilization via damping injection). The resulting controlled robot is passive and very reliable. Classical PD control shows how passivity-based design for fully actuated robot manipulators is a relatively ‘easy’ task.
Although passivity-based control is a well-established control approach, in applications few practical designs go beyond PD control and quadratic potentials. Likewise, established methods such as virtual fixtures [12] and impedance controllers too often remain within the comfort of proportional-derivative reaction to the position error. In general, it is difficult to connect the design of a complex controller with the specific features of the task. In control-theoretic terms, the challenge is finding the relevant task-dependent metric and then optimizing the structure and parameters of the controller for this metric. Picture the design of a controller for a surgical laparoscopic task [13, 14, 15], where the surgical tool is constrained to pass through a small, but somewhat compliant aperture into the patient’s body. Precise position or force control are insufficient to tackle the complex interaction required by the task. Attacking this through the mathematical derivation of a specific potential energy of passivity-based control appears to fall short. In fact, what kind of ‘energy-shaping’ is best for this task? Even when the controlled impedance of the robot is considered [16], the desired impedance is often chosen heuristically [17]. Pragmatically, the review [18] concludes that efforts are needed “(…) to establish more systematic guidance or methodology on how to specify impedance parameters [such as desired inertia, damping, and stiffness] to reflect the basic dynamic interaction requirements of robotic manipulations raised from customers (…)”.
@@
The design of controllers as virtual mechanisms is the philosophy of virtual model control [19, 20]. Conceptualizing the controller as a virtual mechanism falls within the framework of control through interconnection [3, 21, 11, 22], but opens the design space to mechanical intuition and the ability to customize the structure of the controller to the specific features of the task. For the case of laparoscopic surgery, we can design a compliant virtual mechanism that connects to the surgical tool, constraining its interaction with soft tissues. An illustration is provided in Fig. 1. This is neither position nor force control. It shows connections with impedance control, but without enforcing a specific desired impedance. The shaping of the energy is not reduced to the manipulation of mathematical terms, but is determined by the geometry and dynamics of the virtual mechanism and by the virtual interconnection with the surgical tool, both driven by mechanical intuition.
Nonlinearities are often seen as problematic dynamics to be removed. However, in a virtual mechanism, a nonlinear spring serves as a tool to stiffen or saturate at different extensions. Nonlinear mechanical linkages become a tool for shaping the kinetic and potential energy of the system intuitively without designing custom distance functions for energy-shaping potential fields. Virtual mechanisms are easy to implement using virtual operation space forces. Their implementation is easy to debug, as the controller action can be understood in terms of the physics of the virtual mechanism. If the virtual mechanism is passive, then the closed loop is passive, guaranteeing convergence to a minima of the potential energy of the controlled robot. In fact, the virtual mechanism approach provides robustness for the sim-to-real transition: even with geometric/inertial modeling errors, so long as the robot and controller are passive, we have some level of stability guarantee.
@@
For optimal tuning, we are inspired by recent work on scientific-machine learning: the practise of combining scientific models with gradient-based optimization [23]. Recent developments in automatic differentiation of ordinary differential equation (ODE) solvers have opened the door to efficient optimizations involving continuous-time dynamical systems, attracting increasing attention in robotics [24, 25, 26, 27, 28, 29]. Scientific machine learning has been used in related applications in physical system identification [30, 31, 32]; to find optimal state feedback control policies [33]; to perform optimal energy shaping for a pendulum, tuning a controller which is a passive neural network [34]; to design an energy shaping controller to generate desired closed-loop oscillatory trajectories on an undamped double pendulum [35].
@@
In this paper, we embrace this philosophy. We consider the robot as an open system whose input/output behavior must be optimized. The optimization of the controller does not rely on general-purpose architectures, such as neural networks, autoencoders, etc. In contrast, our ‘model’ (in the sense of machine learning) is the virtual mechanism designed by the user. The expressivity of the controller, that is, the ability to generate a wide family of control policies, is regulated by the level of parameterization of the virtual mechanism, ranging from stiffness and damping parameters to length, inertia, and other geometrical features of its mechanical linkages. These parameters are optimized using loss functions that are surrogates of classical performance control metrics, namely the $\mathcal{L}_{2}$ and $\mathcal{L}_{\infty}$ metrics [36]. Compared to [33], the use of virtual mechanisms leads to dynamic feedback policies. Passivity is not enforced through constraints as in [34], but inherited from the structure of the controller.
@@
Computing a gradient through an ODE solver ‘properly’ allows us to compute quickly, thus handle the optimization over the dynamics of the controlled robot efficiently. Using this approach to tune virtual mechanisms can meet the expectations of the review [18] for systematic parameter specification. At the same time, our approach guarantees stability/robustness certificates, which are often missing in state-of-the-art machine learning methods (sim-to-real gap), limiting their applicability in critical domains such as surgical robotics, where a higher level of certainty about controller stability/robustness is required.
@@
The contributions of this paper include: (i) Design of passivity-based controllers as virtual mechanisms, discussed in Sections II and III. These sections extend the early results of [37]. (ii) The adaptation of $\mathcal{L}_{2}$ and $\mathcal{L}_{\infty}$ performance metrics for tractable optimization based on automatic differentiation. This is discussed in Section IV-A. (iii) A tuning approach based on scientific machine learning implemented by algorithmic differentiation of continuous-time simulations, discussed in Section IV-C. This extends the approach of [38], based on linear matrix inequalities [39]. (iv) Simulation results for the 1DOF and 7DOF robot in a surgical setting. We also provide experimental validation, as shown in Section V-B.
II Virtual mechanisms fundamentals
II-A Virtual forces
@@
To design a controller as a virtual mechanism we need a way to realize the virtual forces arising from the interconnection of virtual mechanism and robot manipulator. These virtual forces must be emulated through the available actuation. Thus, for fully actuated robots, the objective is to drive the actuators of the robot to realise the effect that these virtual forces would cause on the dynamics of the robot, using standard methods from Cartesian and operation space control [10], [40], [41]. In what follows, we confine the discussion to virtual forces for reasons of simplicity, but the approach can be extended to include torques.
@@
Consider the robot in Figure 2 with generalized coordinates ${\bm{q}}$. The generic point ${\bm{z}}$ is a point rigidly attached to one of the robot’s links. We want to realize the effect of a virtual force ${\bm{F}}$ at ${\bm{z}}$. The position of ${\bm{z}}$ in some robot frame (e.g. the robot’s base frame) can be represented as a function of the generalized coordinates ${\bm{q}}$, namely ${\bm{z}}=h_{{\bm{z}}}({\bm{q}})$. The velocity reads
@@
where $J_{{\bm{z}}}({\bm{q}})=\frac{\partial h_{{\bm{z}}}({\bm{q}})}{\partial{\bm{q}}}.$ Both ${\bm{F}}$ and ${\bm{z}}$ are vectors of three elements. By the principle of virtual work [40], the joint actuation
@@
The pairs $({\bm{F}},\dot{\bm{z}})$ and $({\bm{u}},\dot{{\bm{q}}})$ are ports [42] and the Jacobian $J_{\bm{z}}({\bm{q}})$ models a lossless interconnection between them. The virtual force ${\bm{F}}$, and its effect on the robot dynamics, are then physically realized via ${\bm{u}}$.
The effect of multiple virtual forces ${\bm{F}}_{i}$ at generic points ${\bm{z}}_{i}$, $i\in\{1,\dots,n\}$, is captured by superposition[16]:
@@
II-B Virtual elements
The virtual mechanism is built of primitive mechanical elements such as ideal springs, dampers, and inerters combined into virtual structures. Figure 3 shows the symbols adopted in this paper. For simplicity we restrict the discussion to one degree-of-freedom joints, but the approach can be easily extended to more general joints.
@@ II-B Virtual elements
An ideal spring is defined by its potential energy function $V:\mathbb{R}^{3}\to\mathbb{R}$, which must be lower bounded for the spring to be passive. Let the displacement ${\bm{z}}={\bm{z}}_{1}-{\bm{z}}_{2}$ be the extension of a virtual spring whose terminals are connected to the points ${\bm{z}}_{1}$ and ${\bm{z}}_{2}$ of the robot. The associated forces, ${\bm{F}}_{1}$ and ${\bm{F}}_{2}$, respectively acting on ${\bm{z}}_{1}$ and ${\bm{z}}_{2}$, are given by the negative gradient of this potential: ${\bm{F}}_{1}^{T}=-\frac{\partial V({\bm{z}})}{\partial{\bm{z}}_{1}}$ and ${\bm{F}}_{2}^{T}=-\frac{\partial V({\bm{z}})}{\partial{\bm{z}}_{2}}$. The virtual force of the spring reads
@@
A linear spring takes $V({\bm{z}})=\frac{1}{2}k{\bm{z}}^{T}{\bm{z}}$, where $k$ is a positive scalar. The corresponding virtual force is ${\bm{F}}=k{\bm{z}}$. Impedance controllers in robotics are often limited to linear springs. However, in applications, it is useful to adopt nonlinear springs, generated by nonquadratic potentials. For example, the potential energy $V({\bm{z}})=\frac{\sigma^{2}}{k}\ln(\cosh(k|{\bm{z}}|/\sigma))$ leads to a virtual force of the form ${\bm{F}}=\sigma\tanh(k|{\bm{z}}|/\sigma)\frac{{\bm{z}}}{|{\bm{z}}|}$. This ‘tanh-spring’ models a useful saturation. In the context of robotic control, this component cannot apply a force greater than $\sigma$, which can be used to build mechanisms with a limited maximum force, for safer human-robot interaction. Similarly, the potential energy $k\frac{({\bm{z}}^{T}{\bm{z}})^{p}}{2p}$, $p>1$, makes the force grow rapidly for $|{\bm{z}}|>1$, which can be used to enforce constraints on robot motion.
@@
The inerter [43] can be considered a generalized inertia/mass. It is used to add inertance to a mechanism. The force of a linear inerter is
@@
As virtual springs and dampers, the action of a virtual inerter is realized by (6) whenever ${\bm{z}}$ is associated to two generic points ${\bm{z}}={\bm{z}}_{1}-{\bm{z}}_{2}$. In what follows, we avoid attaching virtual inerters to the robot directly, as it is difficult to accurately determine the robot accelerations. We connect inerters to points that are internal to the virtual mechanism that characterize the controller, whose accelerations can be computed analytically. An equivalent choice is often made in many impedance controllers, which shape only the stiffness and damping by leaving the robot’s inertia matrix unchanged (the ‘target’ matrix corresponds to the nominal matrix) [44, 45, 17, 46].
@@
where $x$ is an additional ‘state’ variable. This nonlinear inerter is passive if $V$ is lower bounded. In this form, the inerter is an admittance (force as input and motion as output). The analogy with the spring follows by rewriting (5) as
@@
II-C Virtual structures and implementation
Virtual springs, dampers, and inerters are combined into virtual structures defined by a set of rigid bodies connected by rigid and prismatic joints. Extensions to higher-d.o.f. joints are straightforward. Together, these compose the virtual mechanism that represents the controller. The controller has dynamics, since rigid bodies and joints introduce new states. The action of the controller is ultimately determined by the interconnection with the robot, which is always realized through springs and dampers. Figure 5 provides an illustration of the virtual model controller and its interconnection with the robot.
The virtual mechanism is a passive dynamic extension of the robot dynamics (control by interconnection [3]). A virtual mechanism can model mechanical constraints, such as virtual fixtures [12], but its dynamics allow for the modeling of more general dynamical constraints, virtual tools (as shown in Figure 1), and even virtual copies of the robot (digital twin). The geometry of the virtual structure combined with the tuning of the parameters of springs, dampers, and inerters shape the impedance [16] of the interaction port $({\bm{F}}_{e},{\bm{z}}_{e})$, which captures the response of the robot to the external world.
We endow any virtual structure with mass/inertia/inertance such that its inertia matrix is nonsingular and it can be simulated as a second-order-dynamics. The implementation of the controller is captured in Figure 5. Arrows indicate the flow of information. The robot and the virtual structure act as admittances, taking force inputs and returning motion; the interface elements act as impedances taking input from motion and returning the forces. To implement the virtual structures, they are simulated in real time, as
$\ddot{{\bm{q}}}_{c}=M_{c}({\bm{q}}_{c})^{-1}({\bm{u}}_{c})$ (11)
where ${\bm{q}}_{c}$ represents the generalized coordinates of the virtual model controller, and $M_{c}$ is the combined inertia/inertance matrix (the latter due to inerters directly attached to the virtual structure).

### v3

@@ I Introduction
Passivity-based control methods are crucial for designing of stable and reliable robot controllers [Takegaki1981, Ortega2001, Ortega2008, Chopra2022, Secchi2007, Slotine1987, Ortega1989]. Passivity-based approaches consider the energy of the robot-controller interconnection (energy shaping) and regulate its dissipation (damping injection) to ensure stability. The aim is a controlled robot that is passive, thus stable when interacting with other passive devices. The passivity theorem [vanderSchaft2017, Sepulchre1997] guarantees stability even if both the robot model and the environment are uncertain, provided they are passive.
A classical example is proportional derivative (PD) control plus gravity compensation in end-effector space [Khatib1987, Ortega1998]. The proportional action has the mechanical interpretation of a spring. Combined with gravity compensation, it shapes the potential energy of the controlled robot. The derivative action has the mechanical interpretation of a damper, which has the effect of slowing down the robot motion towards the minimum of the shaped potential energy (stabilization via damping injection). The resulting controlled robot is passive and very reliable. In fact, small perturbations or uncertainties in the robot dynamics would not affect the robot motion in a significant way.
Although passivity-based control is a well-established control approach, in applications few practical designs go beyond PD control and quadratic potentials. In fact, related (and powerful!) approaches such as virtual fixtures [Bowyer2014] and impedance controllers too often remain within the comfort of proportional-derivative reaction to the position error. In general, it is difficult to connect the design of a complex controller with the specific features of the task. In control-theoretic terms, the challenge is finding the relevant task-dependent metric and then optimizing the structure and parameters of the controller for this metric. Picture the design of a controller for a surgical laparoscopic task [Su2019, Marinho2019, Su2020a], where the surgical tool is constrained to pass through a small, but somewhat compliant aperture into the patient’s body. Precise position or force control are insufficient to tackle the complex interaction required by the task. Attacking this through the mathematical derivation of a specific potential energy of passivity-based control appears to fall short. In fact, what kind of ‘energy-shaping’ is best for this task? Even when the controlled impedance of the robot is considered [Hogan1984], the desired impedance is often chosen heuristically [Hogan2022]. Pragmatically, the review [Song2019] concludes that efforts are needed “(…) to establish more systematic guidance or methodology on how to specify impedance parameters [such as desired inertia, damping, and stiffness] to reflect the basic dynamic interaction requirements of robotic manipulations raised from customers”.
@@
The design of controllers as virtual mechanisms is the philosophy of virtual model control [Pratt1995b, Joly1995]. Conceptualizing the controller as a virtual mechanism falls within the framework of control through interconnection [Ortega2008, Bloch2000, Ortega1998, Ortega2004], but opens the design space to mechanical intuition and the ability to customize the structure of the controller to the specific features of the task. Examples of virtual model control have been developed in locomotion of bipedal walking robots [Pratt1998, JianjuenHrr1998], exoskeletons [Ekkelenkamp2007], quadrupeds [Desai2014, Chen2020a], virtual fixture approaches [Bowyer2014] and telerobotics [Joly1995, Xu2016]. One of the contributions of this paper is to revisit virtual model control design at fundamental level, starting from the systematic characterization of its basic elements, to arrive to the task-oriented design of passive controllers.
For the case of laparoscopic surgery, we can design a compliant virtual mechanism that connects to the surgical tool, constraining its interaction with soft tissues. An illustration is provided in Fig. 1. This is neither position nor force control. It shows connections with impedance control, but without enforcing a specific desired impedance. The shaping of the energy is not reduced to the manipulation of mathematical terms, but is determined by the geometry and dynamics of the virtual mechanism and by the virtual interconnection with the surgical tool, both driven by mechanical intuition. For the specific case of laparoscopic surgery, our approach shows connections with energy shaping RCM control approaches [Su2020a, Su2020, Su2019, Sadeghian2019, Sandoval2018a, aghakhani_task_2013, Funda1996]. However, our energy-shaping is driven (and constrained) by the geometrical and mechanical representation of a virtual mechanism. This is a qualitative difference which leads to the ability to parametrize entire families of suitable controllers, that can be optimised in a task-oriented way.
Often in robotic control, nonlinearities are seen as problematic dynamics to be removed. In contrast, in a virtual mechanisms controller, a nonlinear spring can stiffen or saturate at different extensions, therefore providing complex controlled feedback forces to serve different performance requirements. Nonlinear mechanical linkages become a tool for shaping the kinetic and potential energy of the system intuitively without designing custom distance functions for energy-shaping potential fields. Virtual mechanisms are easy to implement using virtual operation space forces. Their implementation is easy to debug, as the controller action can be understood in terms of the physics of the virtual mechanism. If the virtual mechanism is passive, then the closed loop is passive, guaranteeing convergence to a minima of the potential energy of the controlled robot. In fact, the virtual mechanism approach provides robustness for the sim-to-real transition: even with geometric/inertial modeling errors, so long as the robot and controller are passive, we have some level of stability guarantee.
@@
For optimal tuning, we are inspired by recent work on scientific-machine learning: the practise of combining scientific models with gradient-based optimization [Rackauckas2021]. Recent developments in automatic differentiation of ordinary differential equation (ODE) solvers have opened the door to efficient optimizations involving continuous-time dynamical systems, attracting increasing attention in robotics [Rackauckas2021a, Kim2021, Chen2019, Howell2023, Giftthaler2017, Giftthaler2018]. Scientific machine learning has been used in related applications in physical system identification [Raissi2019, Greydanus2019, Brunton2019]; to find optimal state feedback control policies [Sandoval2022]; to perform optimal energy shaping for a pendulum, tuning a controller which is a passive neural network [Massaroli2022]; to design an energy shaping controller to generate desired closed-loop oscillatory trajectories on an undamped double pendulum [Wotte2022].
@@
In this paper, we embrace this philosophy. We consider the robot as an open system whose input/output behavior must be optimized. The optimization of the controller does not rely on general-purpose architectures, such as neural networks, autoencoders, etc. In contrast, our ‘model’ (in the sense of machine learning) is the virtual mechanism designed by the user. The expressivity of the controller, that is, the ability to generate a wide family of control policies, is regulated by the level of parameterization of the virtual mechanism, ranging from stiffness and damping parameters to length, inertia, and other geometrical features of its mechanical linkages. These parameters are optimized using loss functions that are surrogates of classical performance control metrics, namely the $\mathcal{L}_{2}$ and $\mathcal{L}_{\infty}$ metrics [Zhou1998]. Compared to [Sandoval2022], the use of virtual mechanisms leads to dynamic feedback policies. Passivity is not enforced through constraints as in [Massaroli2022], but inherited from the structure of the controller.
@@
Computing a gradient through an ODE solver ‘properly’ allows us to compute quickly, thus handle the optimization over the dynamics of the controlled robot efficiently. Using this approach to tune virtual mechanisms can meet the expectations of the review [Song2019] for systematic parameter specification. At the same time, our approach guarantees stability/robustness certificates, which are often missing in state-of-the-art machine learning methods (sim-to-real gap), limiting their applicability in critical domains such as surgical robotics, where a higher level of certainty about controller stability/robustness is required.
@@
The contributions of this paper include: (i) Design of passivity-based controllers as virtual mechanisms, discussed in Sections III and IV. These sections extend the early results of [Larby2022]. (ii) The adaptation of $\mathcal{L}_{2}$ and $\mathcal{L}_{\infty}$ performance metrics for tractable optimization based on automatic differentiation. This is discussed in Section V-A. (iii) A tuning approach based on scientific machine learning implemented by algorithmic differentiation of continuous-time simulations, discussed in Section V-C. This extends the approach of [Larby2022b], based on linear matrix inequalities [Boyd1994]. (iv) Simulation results for the 1DOF and 7DOF robot in a surgical setting. We also provide experimental validation, as shown in Section VI-B.
II Passivity Fundamentals
Even if passivity is widely used and well known in the context robot control [Takegaki1981, Ortega2001, Ortega2008, Chopra2022, Secchi2007, Slotine1987, Ortega1989, vanderSchaft2017, Sepulchre1997], we briefly revisit the main concepts to make this paper self-contained. Mechanical systems without internal sources are passive, that is, any variation of their mechanical energy is bounded by the energy supplied externally. Power may flow into/out of a mechanical system through interconnection ports consisting of a co-located (generalized) force and velocity $(\bm{F},\dot{\bm{z}})$. A mechanical system without internal sources and with a single port $(\bm{F},\dot{\bm{z}})$ satisfies.
$E(t_{1})-E(t_{0})\leq\int_{t_{0}}^{t_{1}}\dot{\bm{z}}^{T}\bm{F}dt,$ (1)
where $E$ represents the mechanical energy of the system, and $t_{0}\leq t_{1}$ are generic time instants. This inequality is the footprint of a passive system. It shows how the variation of mechanical energy is bounded by the supplied energy. Without external supply, the energy of the mechanical system cannot increase (no internal sources). When the energy of the system is differentiable, this is equivalent to
$\dot{E}\leq\dot{\bm{z}}^{T}\bm{F}.$ (2)
The interconnection of passive mechanical system is also passive. To see this, consider two passive mechanical systems with energy $E_{1}$ and $E_{2}$, respectively. The interconnection of these system through two generic ports, $(\bm{F}_{1},\dot{\bm{z}_{1}})$ on the first system and $(\bm{F}_{2},\dot{\bm{z}_{2}})$ on the second system, must satisfy either,
$\dot{\bm{z}}_{1}=\dot{\bm{z}}_{2}=\dot{\bm{z}}\qquad\bm{F}_{1}+\bm{F}_{2}=\bm{F},$ (3a)
or
$\dot{\bm{z}}_{1}+\dot{\bm{z}}_{2}=\dot{\bm{z}}\qquad\bm{F}_{1}=\bm{F}_{2}=\bm{F},$ (3b)
where $(\bm{F},\bm{z})$ are additional co-located force and velocity variables. Then, the interconnected system has total energy
$E=E_{1}+E_{2}$ (4)
and is passive with respect to the ‘new’ port $(\bm{F},\dot{\bm{z}})$. For instance,
$\dot{E}=\dot{E}_{1}+\dot{E}_{2}\leq\dot{\bm{z}}_{1}^{T}\bm{F}_{1}+\dot{\bm{z}}_{2}^{T}\bm{F}_{2}\leq\dot{\bm{z}}^{T}\bm{F}$ (5)
The feature that passivity is preserved by mechanical interconnections is at the core of the use of virtual mechanisms for control design. The idea is that robots are passive and can be controlled by ‘interconnecting’ them to virtual mechanisms, which are also passive. Later, the action of this virtual mechanism is realized through the controlled action of the available motor actuation.
It is a standard result of passive system theory that, in the absence of external forces and in the presence of sufficient damping, all trajectories of the controlled robot converge to some resting state. These correspond to the equilibria of the system, given by zero velocity and by generalized positions that are extremal of the potential energy of the controlled system. This is because without power flowing into the system, $\bm{F}=0\implies\dot{\bm{z}}^{T}\bm{F}=0$, the rate of change of energy is strictly negative [^1] Due to the ‘sufficient’ damping mentioned above unless the system is at rest [Ortega1989, Ortega1998].
III Virtual mechanisms fundamentals
III-A Virtual forces
@@
To design a controller as a virtual mechanism we need a way to realize the virtual forces arising from the interconnection of virtual mechanism and robot manipulator. These virtual forces must be emulated through the available actuation. Thus, for fully actuated robots, the objective is to drive the actuators of the robot to realise the effect that these virtual forces would cause on the dynamics of the robot, using standard methods from Cartesian and operation space control [Khatib1987], [Spong1989], [Siciliano1999]. In what follows, we confine the discussion to virtual forces for reasons of simplicity, but the approach can be extended to include torques.
@@
Consider the robot in Fig. 2 with generalized coordinates $\bm{q}$}. The generic point $\bm{z}$} is a point rigidly attached to one of the robot’s links. We want to realize the effect of a virtual force $\bm{F}$} at $\bm{z}$}. The position of $\bm{z}$} in some robot frame (e.g. the robot’s base frame) can be represented as a function of the generalized coordinates $\bm{q}$}, namely ${\bm{z}}=h_{{\bm{z}}}({\bm{q}})$. The velocity reads
@@
where $J_{{\bm{z}}}({\bm{q}})=\frac{\partial h_{{\bm{z}}}({\bm{q}})}{\partial{\bm{q}}}.$ Both $\bm{F}$} and $\bm{z}$} are vectors of three elements. By the principle of virtual work [Spong1989], the joint actuation
@@
The pairs $({\bm{F}},\dot{\bm{z}})$ and $({\bm{u}},\dot{{\bm{q}}})$ are ports [vanderSchaft2020] and the Jacobian $J_{\bm{z}}({\bm{q}})$ models a lossless interconnection between them. The virtual force $\bm{F}$}, and its effect on the robot dynamics, are then physically realized via $\bm{u}$}.
The effect of multiple virtual forces ${\bm{F}}_{i}$ at generic points ${\bm{z}}_{i}$, $i\in\{1,\dots,n\}$, is captured by superposition[Hogan1984]:
@@
III-B Virtual elements
The virtual mechanism is built of primitive mechanical elements such as ideal springs, dampers, and inerters combined into virtual structures. Fig. 3 shows the symbols adopted in this paper. For simplicity we restrict the discussion to one degree-of-freedom joints, but the approach can be easily extended to more general joints.
@@ III-B Virtual elements
An ideal spring is defined by its potential energy function $V:\mathbb{R}^{3}\to\mathbb{R}$, which must be lower bounded for the spring to be passive. Let the displacement ${\bm{z}}={\bm{z}}_{1}-{\bm{z}}_{2}$ be the extension of a virtual spring whose terminals are connected to the points ${\bm{z}}_{1}$ and ${\bm{z}}_{2}$ of the robot. The external forces, ${\bm{F}}_{1}$ and ${\bm{F}}_{2}$, respectively acting on ${\bm{z}}_{1}$ and ${\bm{z}}_{2}$, balance the force in the spring, and are given by the gradient of this potential: ${\bm{F}}_{1}^{T}=\frac{\partial V({\bm{z}})}{\partial{\bm{z}}_{1}}$ and ${\bm{F}}_{2}^{T}=\frac{\partial V({\bm{z}})}{\partial{\bm{z}}_{2}}$. The virtual force of the spring reads
@@
Any such spring is passive: taking the derivative of $V$,
$\dot{V}$ $=\dot{\bm{z}}_{1}^{T}\frac{\partial V}{\partial\bm{z}_{1}}+\dot{\bm{z}}_{2}^{T}\frac{\partial V}{\partial\bm{z}_{2}}$ (12a)
$=\dot{\bm{z}}_{1}^{T}\bm{F}_{1}+\dot{\bm{z}}_{2}^{T}\bm{F}_{2},$ (12b)
we show that the spring satisfies (2). Note the equality sign: the spring stores energy but does not dissipate energy.
A linear spring takes $V({\bm{z}})=\frac{1}{2}k{\bm{z}}^{T}{\bm{z}}$, where $k$ is a positive scalar. The corresponding virtual force is ${\bm{F}}=k{\bm{z}}$. Impedance controllers in robotics are often limited to linear springs. However, in applications, it is useful to adopt nonlinear springs, generated by nonquadratic potentials. For example, the potential energy $V({\bm{z}})=\frac{\sigma^{2}}{k}\ln(\cosh(k|{\bm{z}}|/\sigma))$ leads to a virtual force of the form ${\bm{F}}=\sigma\tanh(k|{\bm{z}}|/\sigma)\frac{{\bm{z}}}{|{\bm{z}}|}$, as plotted in Fig. 4. This ‘tanh-spring’ models a useful saturation. In the context of robotic control, this component cannot apply a force greater than $\sigma$, which can be used to build mechanisms with a limited maximum force, for safer human-robot interaction. Similarly, the potential energy $k\frac{({\bm{z}}^{T}{\bm{z}})^{p}}{2p}$, $p>1$, makes the force grow rapidly for $|{\bm{z}}|>1$, which can be used to enforce constraints on robot motion.
@@
The inerter [Smith2002] can be considered a generalized inertia/mass. It is used to add inertance to a mechanism. The force of a linear inerter is
@@
As virtual springs and dampers, the action of a virtual inerter is realized by (11) whenever $\bm{z}$} is associated to two generic points ${\bm{z}}={\bm{z}}_{1}-{\bm{z}}_{2}$. In what follows, we avoid attaching virtual inerters to the robot directly, as it is difficult to accurately determine the robot accelerations. We connect inerters to points that are internal to the virtual mechanism that characterize the controller, whose accelerations can be computed analytically. An equivalent choice is often made in many impedance controllers, which shape only the stiffness and damping by leaving the robot’s inertia matrix unchanged (the ‘target’ matrix corresponds to the nominal matrix) [Averta2020, Lachner2022, Hogan2022, Lawrence1988].
@@
where $x$ is an additional ‘state’ variable, a generalized momentum. This nonlinear inerter is passive if $V$ is lower bounded. In this form, the inerter is an admittance (force as input and motion as output). The analogy with the spring follows by rewriting (10) as
@@
III-C Virtual structures and implementation
Virtual springs, dampers, and inerters are combined into virtual structures defined by a set of rigid bodies connected by rigid and prismatic joints. Extensions to higher-d.o.f. joints are straightforward. Together, these compose the virtual mechanism that represents the controller. The controller has dynamics, since rigid bodies and joints introduce new states. The action of the controller is ultimately determined by the interconnection with the robot, which is always realized through springs and dampers. Fig. 5 provides an illustration of the virtual model controller and its interconnection with the robot.
The virtual mechanism is a passive dynamic extension of the robot dynamics (control by interconnection [Ortega2008]). A virtual mechanism can model mechanical constraints, such as virtual fixtures [Bowyer2014], but its dynamics allow for the modeling of more general dynamical constraints, virtual tools (as shown in Fig. 1), and even virtual copies of the robot (digital twin). The geometry of the virtual structure combined with the tuning of the parameters of springs, dampers, and inerters shape the impedance [Hogan1984] of the interaction port $({\bm{F}}_{e},{\bm{z}}_{e})$, which captures the response of the robot to the external world.
We endow any virtual structure with mass/inertia/inertance such that its inertia matrix is nonsingular and it can be simulated as a second-order-dynamics. The implementation of the controller is captured in Fig. 5. Arrows indicate the flow of information. The robot and the virtual structure act as admittances, taking force inputs and returning motion; the interface elements act as impedances taking input from motion and returning the forces [^2] Note that this uses the force-effort analogy, while there are compelling reasons to use a force-flow analogy, for a description see [Smith2020, Section 3]. To implement the virtual structures, they are simulated in real time, as
$M_{c}({\bm{q}}_{c})\ddot{{\bm{q}}}_{c}+C_{c}({\bm{q}}_{c},\dot{{\bm{q}}}_{c})\dot{{\bm{q}}}_{c}={\bm{u}}_{c}$ (17)
where ${\bm{q}}_{c}$ represents the generalized coordinates of the virtual model controller, $M_{c}$ is the combined inertia/inertance matrix (the latter due to inerters directly attached to the virtual structure), and $C_{c}({\bm{q}}_{c},\dot{{\bm{q}}}_{c})$ models Coriolis/gyroscopic effects.